OpenStudy (anonymous):

Write an equation in slope-intercept form for the line that passes through (-5,3) and is perpendicular to the line described by y= 5x.

Parth (parthkohli):

Okay, so do you know what the slope of a line perpendicular to another is?

OpenStudy (anonymous):

uhuhuhuh no not really

OpenStudy (anonymous):

Perpendicular Means Product of their Slopes = -1 ie m1*m2 = -1

Parth (parthkohli):

It's like the opposite of the reciprocal. Got it? So if the slope is 8, then the slope of the perpendicular line is: 1) first find the reciprocal. 1/8 2) then the opposite -1/8

Parth (parthkohli):

So, can you do what is shown in the above?

OpenStudy (anonymous):

y= 5x y = mx + c m = slope So in y= 5x m = 5 5 * m2 = -1 m2 = -1/5 nw use point slope form y-y1=m(x-x1) using (-5,3) m = -1/5

Parth (parthkohli):

Instead, do it for 5. What is the opposite-reciprocal of 5?

Parth (parthkohli):

How about we have an interaction with the user who asked it?

OpenStudy (anonymous):

@vanessaloar u got it ????

OpenStudy (anonymous):

yess, I do I just couldn't remember how to do this and Im on my final exam on my online school thanks guys:)

Parth (parthkohli):

You're welcome ;)

OpenStudy (anonymous):

Welcome..)

OpenStudy (anonymous):

<3 :)